OpenStudy (anonymous):

I'm trying to evaluate the limit of a function but I keep getting the wrong answer. What am I doing wrong? Equation coming in next post.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1}\frac{ 2x^2+x-3 }{ x^2-x }\]

OpenStudy (whpalmer4):

How are you working it, and what do you get?

hartnn (hartnn):

can you factor the numerator ?

OpenStudy (anonymous):

\[\frac{ (2x+3)(x-1) }{ x(x-1) }\]

OpenStudy (anonymous):

I have factored the denominator to (x-1)(x+3/2) so I get \[\lim_{x \rightarrow 1}\frac{ x+\frac{ 3 }{ 2 } }{ x }\] in the end and get the result 5/2 which is wrong o.o

OpenStudy (anonymous):

\[\huge \frac{ x+\frac{ 3 }{ 2 } }{ \frac{ x }{ 2} }\]

hartnn (hartnn):

numerator factors to 2 (x-1)(x+3/2)

OpenStudy (anonymous):

Omg, just saw that 2 >.< Thanks so much, guys!!!